OpenStudy (swagster23780):

Which ratio is equivalent to 18/24 ? A.5/8 B.2/3 C.10/12 D.3/4

OpenStudy (abdullahsadiqi):

do Make the 18/24 simple to give you one of the answer below

OpenStudy (abdullahsadiqi):

\[\frac{ 18 }{24 } =\frac{ 9 }{ 12}=\frac{ 3 }{ 4 }\]

OpenStudy (abdullahsadiqi):

in rules of ratio there should be mentioned that we can divide the first and second number of a ratio by same number and there in ratio there will be no difference .

OpenStudy (abdullahm):

Divider the numerator and denominator by 3 to get your answer. 18 / 3 = ? 24 / 3 = ?

OpenStudy (abdullahsadiqi):

do scaling down !